Calculate the numbers of formula units of 50.0 mol NaNO3

1
Expert's answer
2021-04-16T05:09:32-0400

"\frac{m}{M} = \frac{N}{Na}"

"N = 50\times 6.02 \times 10^{23} = 3.01\times 10^{25}"

Answer:

"N = 3.01 \times 10^{25}"